a body of velocity-2 kmph strikes ground from height h .another body of same mass from same height with initial velocity 4 kmph is thrown down .final v of 2nd body?

Grade:12

1 Answers

Ritika Bagokar
7 years ago
h=2t+1/2gt^2      &   h=4t+1/2t^2   equate        
the two equations   to get    t=2sec 
put in either of equations to get h   
ie h=16   
use v^2=u^2+2gh
v=√4^2 +2g*16 =√336=18.33
hence ans i 18.33